# GeoZero: Incentivizing Reasoning from Scratch on Geospatial Scenes

### v2

@@ Abstract
Multimodal large language models (MLLMs) have undergone rapid development in advancing geospatial scene understanding. Recent studies have sought to enhance the reasoning capabilities of remote sensing MLLMs, typically through cold-start training with elaborately curated chain-of-thought (CoT) data. However, this approach not only incurs substantial annotation costs but also introduces human biases that may limit the diversity of model reasoning. To address these challenges, we propose GeoZero, a framework that enables MLLMs to perform geospatial reasoning without any predefined CoT supervision. Specifically, we construct two datasets, GeoZero-Instruct and GeoZero-Hard. GeoZero-Instruct allows the model to acquire preliminary geospatial knowledge through supervised fine-tuning, while GeoZero-Hard stimulates deep reasoning during the subsequent reinforcement learning stage. Furthermore, we introduce Answer-Anchored Group Relative Policy Optimization (A2GRPO), where the reasoning process is regularized by the model’s own answers, encouraging diverse yet accurate thinking. Extensive experiments on multiple remote sensing vision–language benchmarks demonstrate that GeoZero not only surpasses existing state-of-the-art methods but also fosters universal emergent reasoning capabilities across diverse geospatial tasks. Code, data, and models will be publicly available at GeoZero.
[…]
1 Introduction
@@
2 Related Work
2.1 MLLMs on Geospatial Scenes
@@
2.2 Large Multimodal Reasoning Models
@@ 3 Method
3.1 Overall Training Paradigm
@@
Table 1: Composition of the GeoZero-Raw dataset. “Others” include counting, honesty evaluation, and multi-turn dialogue samples without explicit task identifiers.
@@ 3.2 Dataset Construction
3.2.1 GeoZero-Raw
We mainly focus on four representative remote sensing vision–language tasks: scene classification (SC), visual grounding (VG), visual question answering (VQA), and image captioning (IC). For each task, we collect samples from multiple classical datasets. For instance, the SC task includes data from the training sets of AID [61], RESISC-45 [8], and NASC-TG2 [75], among others. The resulting GeoZero-Raw dataset spans diverse resolutions, sensor modalities, and scene categories, which helps enhance the generalization ability of MLLMs for remote sensing understanding. A detailed composition of the dataset is summarized in Table 1.
@@ 3.2.1 GeoZero-Raw
3.2.2 GeoZero-Instruct and GeoZero-Hard
@@
Notably, to avoid potential image overlap within GeoZero-Raw (as discussed in Section-3.2.1), we remove all samples sharing the same image with GeoZero-Hard. The remaining samples form GeoZero-Instruct. As a result, GeoZero-Hard and GeoZero-Instruct contain approximately 20K and 610K samples, respectively.
3.3 Answer-Anchored Group Relative Policy Optimization
@@ 4.1 Experimental Settings
4.2 Comparison to State-of-the-art MLLMs
Table 2: Performance comparison (%) of different MLLMs on scene classification tasks. The best and second-best results are highlighted in bold and blue, respectively.
@@ 4.2 Comparison to State-of-the-art MLLMs
Table 3: Performance comparison (%) of different MLLMs on visual grounding tasks. The best and second-best results are highlighted in bold and blue, respectively.
@@
Table 4: Performance comparison (%) of different MLLMs on visual question answering tasks. The best and second-best results are highlighted in bold and blue, respectively.
@@
Table 5: Performance comparison (%) of different MLLMs on image captioning tasks. The best and second-best results are highlighted in bold and blue, respectively.
@@
Typical Remote Sensing Vision-Language Tasks. Tables 2-5 summarize the performance of GeoZero against representative state-of-the-art remote sensing MLLMs on four typical tasks: SC, VG, VQA, and IC. For evaluation, SC and VQA are measured by accuracy. VG adopts Acc@0.5, where a prediction is considered correct if the IoU between the predicted box and the ground truth exceeds 0.5. For IC, we report standard captioning metrics, including BLEU-4, CIDEr, and METEOR.
For SC, we evaluate on two widely used benchmarks, UCM [65] and AID [61], following common evaluation practices in recent works [47, 27, 45]. As shown in Table 2, GeoZero achieves superior performance on both datasets, outperforming both general MLLMs and recent remote sensing MLLMs. Notably, GeoZero yields a clear margin on UCM, indicating strong generalization to unseen scenes.
For VG, we report results on RSVG [55] and DIOR-RSVG [67]. As shown in Table 3, GeoZero delivers the best performance on RSVG and remains highly competitive on DIOR-RSVG, approaching the specialized grounding model GeoGround [74]. These results demonstrate that our method effectively strengthens spatial localization and grounding capability.
For VQA, we adopt RSVQA-HR [39] for evaluation and report accuracy on its two question types (Presence and Compare). The results in Table 4 show that GeoZero attains the highest accuracy on both question types, improving substantially over the general-purpose base model while remaining competitive against prior remote sensing MLLMs.
Finally, for IC, we evaluate on RSICD [40]. As shown in Table 5, GeoZero achieves the best CIDEr and METEOR among the compared methods, despite CIDEr not being explicitly optimized in the task-specific answer reward. This suggests that the proposed training framework can improve the overall quality of generated descriptions beyond the directly optimized components.
4.3 Ablation Study
@@
Table 6: Ablation on different training settings. $\bullet$ and $\circ$ indicate whether thinking texts are generated during inference ($\bullet$: with reasoning, $\circ$: without reasoning).
@@
Table 7: Ablation study on the dataset used during RL. “Random” means the data are randomly selected from GeoZero-Raw-val (see Section-D in the supplementary material), while the construction method of the SFT dataset is the same as GeoZero-Instruct. $\bullet$ and $\circ$ follow the same definitions as in Table 6.
@@
Table 8: Ablation on different components of A2GRPO. $\bullet$ and $\circ$ follow the same definitions as in Table 6.
@@
Training Paradigm. We first examine the impact of different training strategies, as summarized in Table 6. Direct RL training (RL only) can trigger reasoning generation ($\bullet$). However, without SFT, the model lacks fundamental geospatial knowledge and performs poorly on challenging benchmarks (e.g., RSVG). Although RFT can partially recover accuracy, it still falls short of the SFT-RL pipeline trained with GeoZero-Hard. In contrast, SFT-only and SFT+RFT achieve strong accuracy but fail to activate reasoning ($\circ$), likely due to the absence of CoT supervision in SFT and the limited scale and diversity of downstream training data (compared with GeoZero-Hard) in RFT, which provide insufficient optimization signals to induce reasoning. By incorporating the constructed hard-sample pool, GeoZero-Hard, into the SFT-RL paradigm (GeoZero w/o RFT), the model maintains high accuracy while consistently activating reasoning, indicating its ability to think from scratch. Further applying dataset-specific RFT (GeoZero w/ RFT) improves both task accuracy and reasoning consistency, validating the effectiveness of our training paradigm.
@@
GeoZero-Hard. We evaluate the impact of the constructed datasets, as shown in Table 7. It can be observed that when RL training uses randomly selected data instead of hard samples, although the model performs well on the UCM dataset, it fails to generate reasoning texts, and the performance significantly decreases on the RSVG dataset. This suggests that the model struggles with reasoning on more challenging scenes when trained with random samples. After RFT, the model reactivates reasoning on the more challenging RSVG dataset, yet its performance remains to be further improved. These results highlight the importance of using hard samples during RL training to help the model engage in genuine reasoning from scratch.
@@
A2GRPO. We further evaluate the effectiveness of the proposed A2GRPO algorithm by examining the influence of the AMTR and the thinking mask. As shown in Table 8, without answer modulation (i.e., $r_{t}=s_{t}$), the emergence of reasoning becomes independent of the answer reward and is further constrained by factors such as KL penalties and redundancy suppression. In this case, without CoT-based cold-start training, the model tends to converge to a risk-averse strategy: simply outputting answers without generating reasoning, so as to maintain stability. When AMTR is applied but the thinking mask is removed, additional constraints hinder the diversity of reasoning generation, thereby degrading model performance, particularly in reasoning-intensive scenarios (e.g., the RSVG dataset). In contrast, A2GRPO encourages constructive reasoning while removing unintended constraints on the thinking process, thereby expanding the reasoning exploration space. As a result, our method effectively activates reasoning while maintaining high answer accuracy, ultimately enhancing the model’s emergent reasoning capability.
@@
In summary, the results in Tables 68 demonstrate that, by leveraging hard samples and proper RL strategies (i.e., AMTR and the thinking mask), we successfully trigger the emergence of thinking in MLLMs on geospatial scenes, a capability that does not appear after SFT.
4.4 Visualization
@@
Figure 3: Relationship between thinking and answering, where the statistics are computed from GeoZero’s prediction results on the DIOR-RSVG test set. (Best view in zoom.)
@@
Thinking v.s. Answering. To further investigate how reasoning relates to final predictions, we analyze the relationship between answer accuracy and both reasoning length and reasoning quality score $s_{t}$. Specifically, all samples are grouped based on their reasoning length and quality score, and the average accuracy within each group is computed. As shown in Figure 3(a), accuracy increases with longer reasoning chains, indicating that extended reasoning generally benefits answer correctness. However, when the reasoning becomes excessively long, i.e., overthinking, the performance slightly drops. Similarly, Figure 3(b) reveals a positive correlation between $s_{t}$ and answer accuracy, indicating that samples with higher reasoning quality scores tend to achieve higher accuracy. Overall, these results suggest a strong association between reasoning quality and task performance, supporting our design choice of encouraging explicit reasoning during RL. More fine-grained analyses are provided in Section-I of the supplementary material.
@@ Appendix B Can General MLLMs Perform Geospatial Reasoning?
To measure whether the model performs explicit reasoning, we adopt a simple yet permissive criterion: a sample is considered as “reasoned” if there exist any texts appearing before the “<answer>” tag in the model’s output. Based on this rule, we calculate the proportion of samples exhibiting reasoning behavior (referred to as the ”thinking activation rate”), and summarize the results in Table 9.
Table 9: Accuracy and thinking activation rate (%) of different models on the UCM and RSVG datasets. GeoZero uses the system prompt (SP) by default. w/SP: with the system prompt (Type1) shown in Figure 8. Acc.: accuracy. TAR: thinking activation rate.
@@ Appendix C Instruction Formatting of GeoZero-Raw
We apply instruction formatting to the GeoZero-Raw dataset. To distinguish different tasks, following the conventions of existing remote sensing MLLMs [28, 68], we prepend task descriptors to user instructions, where [cls], [grounding], [vqa], and [caption] represent scene classification (SC), visual grounding (VG), visual question answering (VQA), and image captioning (IC) tasks, respectively. For samples that do not belong to these four vision–language tasks, no task descriptors are added. Furthermore, to help general-purpose MLLMs better understand the expected output format for different remote sensing vision–language tasks, we introduce task-specific textual hints in addition to the original user instructions.. For example, for VG tasks, the instruction is reformulated as: “Please provide the location of the object in the image: {original grounding text}.” To prevent the model from over-relying on these textual hints, we randomly decide whether to include them with a probability of 50%. In addition, to further enhance model generalization, we design more than 20 variations of textual hints for each task, from which one is randomly selected for each sample. The textual hints for different tasks are shown in Figure 11-14, while the converted samples are presented in Figure 15.
@@ Appendix G Task-Specific Answer Reward $r_{a}$
Table 10: Training hyperparameters for evaluating on different datasets used in scene classification, visual grounding, visual question answering, image captioning tasks and comprehensive benchmarks with GeoZero (w/ RFT).
@@ Appendix H Reinforcement Fine-tuning Settings
The hyperparameters used for reinforcement fine-tuning on different datasets are summarized in Table 10. All other settings follow the reinforcement learning stage described in the main paper.
@@ Appendix I Further Investigation of Reasoning Behavior
I.1 Overall Characteristics of Model Reasoning.
Figure 9(a)–(f) illustrates the distribution of average sample accuracy across different reasoning-related metrics. It should be noted that in the main paper, we conducted a preliminary analysis using two variables: the word-level reasoning length and the thinking-quality score $s_{t}$ (see (a) and (f)). Here, we take a closer look by further decomposing $s_{t}$ into its internal factors. Specifically, we additionally analyze four variables: the normalized length score $l_{s}(T)$, the non-redundancy ratio 1d(T)1-d(T) (which serves as a redundancy penalty in $q_{t}$), the structural thinking score $q_{t}$, and the semantic diversity bonus $b_{d}$. Furthermore, we examine how the sample frequency is distributed across different value intervals of each metric (see (g)–(l)), which helps reveal the overall distribution of reasoning characteristics within the test data. Next, we provide a detailed analysis of each metric individually.
@@
    Reasoning Length. Figure 9(b) shows a positive correlation between $l_{s}(T)$ and accuracy: samples with more appropriate reasoning lengths tend to yield more reliable predictions, consistent with the findings in the main paper. Encouragingly, we find that for the vast majority of samples, the predicted reasoning lengths fall within the expected range, as shown in Figure 9(h), where more than 80% of the samples achieve very high $l_{s}(T)$ values (greater than 0.8). Figure 9(g) further reveals that most reasoning chains have a length between 40 and 80 words, which precisely corresponds to the full-score interval defined in Equation (8).
@@
    Non-Redundancy. As shown in Figures 9(c) and (i), almost no accuracy bars appear when $1-d(T)\leq 0.5$, while the vast majority of samples have values above 0.85. This suggests that the model’s reasoning chains exhibit strong non-redundancy across most samples. However, Figure 9(c) also indicates that variations in reasoning redundancy have negligible influence on model performance.
    Figure 9: The distribution of average accuracy and sample frequency on thinking-related metrics, where the statistics are computed from GeoZero’s predictions on the 7500 samples from the DIOR-RSVG [67] test set. The accuracy of samples are represented by IoU score.
    Figure 10: The distribution of thinking-related variables with respect to accuracy, where the statistics are computed from GeoZero’s predictions on 7,500 samples from the DIOR-RSVG [67] test set. The sample accuracy is represented by the IoU score. The intervals are determined based on the quantiles of the IoU distribution, resulting in non-uniform spacing between the bars.
@@
    Structural Thinking. Figure 9(d) clearly demonstrates that samples with low structural thinking scores $q_{t}$ tend to have low average accuracy, whereas those with higher $q_{t}$ consistently achieve better performance. This indicates that a well-structured reasoning process is a fundamental requirement for maintaining reliable predictions. However, the figure also shows that further increasing $q_{t}$ does not lead to substantial performance gains, implying that, at this stage, the semantic content of reasoning (such as semantic diversity; see (e)) may play a more dominant role. Moreover, Figure 9(j) shows that GeoZero produces well-structured reasoning chains for the majority of samples, reflecting its coherent reasoning capability.
@@
    Semantic Diversity. Figure 9(e) further examines the semantic diversity bonus $b_{d}$. We observe a non-monotonic relationship: moderate semantic diversity correlates with improved accuracy, as it reflects coherent multi-step reasoning. Once $b_{d}$ becomes excessively large, however, the reasoning often degenerates into fragmented or even contradictory statements, suggesting that the model is no longer performing genuine reasoning, which in turn hurts accuracy. Fortunately, Figure 9(k) shows that for the vast majority of samples, the generated reasoning chains have $b_{d}$ values below 0.45, where semantic diversity plays a positive and constructive role.
@@
    Overall Reasoning Quality. Finally, Figures 9(f) and (l) jointly indicate that GeoZero produces reasoning chains with high thinking-quality scores ($s_{t}$) for the majority of samples, which underpins the model’s overall performance across the test set. We also observe that some samples achieve good average accuracy despite relatively low $s_{t}$ values. This phenomenon corresponds to the distinct leftmost bar in Figure 9(j), where the structural thinking scores ($q_{t}$) are also low. This may be attributed to easier cases in which the model can make correct predictions without maintaining a highly organized reasoning structure, e.g., by exhibiting mild redundancy in expression, as seen in the third-to-last bar of Figure 9(i).
@@
I.2 Reasoning Behavior across Samples.
[…]
@@
Interestingly, for the easiest samples (IoU $\in$ [0.95, 1]), the model tends to produce even longer reasoning chains. This leads to mild redundancy, which slightly affects the reasoning-chain quality score $s_{t}$ through the structural thinking score $q_{t}$, as reflected by the lower values of the leftmost bars in Figure 10(c), (d), and (f). However, since these samples are inherently simple, such redundancy does not compromise the model’s prediction accuracy.

### v3

@@ Abstract
Multimodal large language models (MLLMs) have undergone rapid development in advancing geospatial scene understanding. Recent studies have sought to enhance the reasoning capabilities of remote sensing MLLMs, typically through cold-start training with elaborately curated chain-of-thought (CoT) data. However, this approach not only incurs substantial annotation costs but also introduces human biases that may limit the diversity of model reasoning. To address these challenges, we propose GeoZero, a framework that enables MLLMs to perform geospatial reasoning without any predefined CoT supervision. Specifically, we construct two datasets, GeoZero-Instruct and GeoZero-Hard. GeoZero-Instruct allows the model to acquire preliminary geospatial knowledge through supervised fine-tuning, while GeoZero-Hard stimulates deep reasoning during the subsequent reinforcement learning stage. Furthermore, we introduce Answer-Anchored Group Relative Policy Optimization (A2GRPO), where the reasoning process is regularized by the model’s own answers, encouraging diverse yet accurate thinking. Extensive experiments on multiple remote sensing vision–language benchmarks demonstrate that GeoZero not only surpasses existing state-of-the-art methods but also fosters universal emergent reasoning capabilities across diverse geospatial tasks. Code, data, and models are available at GeoZero.
Index Terms:
Remote sensing, multimodal large language models, geospatial reasoning, reinforcement learning
I Introduction
@@
II Related Work
II-A MLLMs on Geospatial Scenes
@@
II-B Large Multimodal Reasoning Models
@@ III Method
III-A Overall Training Paradigm
@@
TABLE I: Composition of the GeoZero-Raw dataset. “Others” include counting, honesty evaluation, and multi-turn dialogue samples without explicit task identifiers.
@@ III-B Dataset Construction
III-B1 GeoZero-Raw
We mainly focus on four representative remote sensing vision–language tasks: scene classification (SC), visual grounding (VG), visual question answering (VQA), and image captioning (IC). For each task, we collect samples from multiple classical datasets. For instance, the SC task includes data from the training sets of AID [61], RESISC-45 [8], and NASC-TG2 [75], among others. The resulting GeoZero-Raw dataset spans diverse resolutions, sensor modalities, and scene categories, which helps enhance the generalization ability of MLLMs for remote sensing understanding. A detailed composition of the dataset is summarized in Table I.
@@ III-B1 GeoZero-Raw
III-B2 GeoZero-Instruct and GeoZero-Hard
@@
Notably, to avoid potential image overlap within GeoZero-Raw (as discussed in Section-III-B1), we remove all samples sharing the same image with GeoZero-Hard. The remaining samples form GeoZero-Instruct. As a result, GeoZero-Hard and GeoZero-Instruct contain approximately 20K and 610K samples, respectively.
III-C Answer-Anchored Group Relative Policy Optimization
@@ IV-A Experimental Settings
IV-B Comparison to State-of-the-art MLLMs
TABLE II: Performance comparison (%) of different MLLMs on scene classification tasks. The best and second-best results are highlighted in bold and blue, respectively.
@@ IV-B Comparison to State-of-the-art MLLMs
TABLE III: Performance comparison (%) of different MLLMs on visual grounding tasks. The best and second-best results are highlighted in bold and blue, respectively.
@@
TABLE IV: Performance comparison (%) of different MLLMs on visual question answering tasks. The best and second-best results are highlighted in bold and blue, respectively.
@@
TABLE V: Performance comparison (%) of different MLLMs on image captioning tasks. The best and second-best results are highlighted in bold and blue, respectively.
@@
Typical Remote Sensing Vision-Language Tasks. Tables II-V summarize the performance of GeoZero against representative state-of-the-art remote sensing MLLMs on four typical tasks: SC, VG, VQA, and IC. For evaluation, SC and VQA are measured by accuracy. VG adopts Acc@0.5, where a prediction is considered correct if the IoU between the predicted box and the ground truth exceeds 0.5. For IC, we report standard captioning metrics, including BLEU-4, CIDEr, and METEOR.
For SC, we evaluate on two widely used benchmarks, UCM [65] and AID [61], following common evaluation practices in recent works [47, 27, 45]. As shown in Table II, GeoZero achieves superior performance on both datasets, outperforming both general MLLMs and recent remote sensing MLLMs. Notably, GeoZero yields a clear margin on UCM, indicating strong generalization to unseen scenes.
For VG, we report results on RSVG [55] and DIOR-RSVG [67]. As shown in Table III, GeoZero delivers the best performance on RSVG and remains highly competitive on DIOR-RSVG, approaching the specialized grounding model GeoGround [74]. These results demonstrate that our method effectively strengthens spatial localization and grounding capability.
For VQA, we adopt RSVQA-HR [39] for evaluation and report accuracy on its two question types (Presence and Compare). The results in Table IV show that GeoZero attains the highest accuracy on both question types, improving substantially over the general-purpose base model while remaining competitive against prior remote sensing MLLMs.
Finally, for IC, we evaluate on RSICD [40]. As shown in Table V, GeoZero achieves the best CIDEr and METEOR among the compared methods, despite CIDEr not being explicitly optimized in the task-specific answer reward. This suggests that the proposed training framework can improve the overall quality of generated descriptions beyond the directly optimized components.
IV-C Ablation Study
@@
TABLE VI: Ablation on different training settings. $\bullet$ and $\circ$ indicate whether thinking texts are generated during inference ($\bullet$: with reasoning, $\circ$: without reasoning).
@@
TABLE VII: Ablation study on the dataset used during RL. “Random” means the data are randomly selected from GeoZero-Raw-val (see Section-D in the supplementary material), while the construction method of the SFT dataset is the same as GeoZero-Instruct. $\bullet$ and $\circ$ follow the same definitions as in Table VI.
@@
TABLE VIII: Ablation on different components of A2GRPO. $\bullet$ and $\circ$ follow the same definitions as in Table VI.
@@
Training Paradigm. We first examine the impact of different training strategies, as summarized in Table VI. Direct RL training (RL only) can trigger reasoning generation ($\bullet$). However, without SFT, the model lacks fundamental geospatial knowledge and performs poorly on challenging benchmarks (e.g., RSVG). Although RFT can partially recover accuracy, it still falls short of the SFT-RL pipeline trained with GeoZero-Hard. In contrast, SFT-only and SFT+RFT achieve strong accuracy but fail to activate reasoning ($\circ$), likely due to the absence of CoT supervision in SFT and the limited scale and diversity of downstream training data (compared with GeoZero-Hard) in RFT, which provide insufficient optimization signals to induce reasoning. By incorporating the constructed hard-sample pool, GeoZero-Hard, into the SFT-RL paradigm (GeoZero w/o RFT), the model maintains high accuracy while consistently activating reasoning, indicating its ability to think from scratch. Further applying dataset-specific RFT (GeoZero w/ RFT) improves both task accuracy and reasoning consistency, validating the effectiveness of our training paradigm.
@@
GeoZero-Hard. We evaluate the impact of the constructed datasets, as shown in Table VII. It can be observed that when RL training uses randomly selected data instead of hard samples, although the model performs well on the UCM dataset, it fails to generate reasoning texts, and the performance significantly decreases on the RSVG dataset. This suggests that the model struggles with reasoning on more challenging scenes when trained with random samples. After RFT, the model reactivates reasoning on the more challenging RSVG dataset, yet its performance remains to be further improved. These results highlight the importance of using hard samples during RL training to help the model engage in genuine reasoning from scratch.
@@
A2GRPO. We further evaluate the effectiveness of the proposed A2GRPO algorithm by examining the influence of the AMTR and the thinking mask. As shown in Table VIII, without answer modulation (i.e., $r_{t}=s_{t}$), the emergence of reasoning becomes independent of the answer reward and is further constrained by factors such as KL penalties and redundancy suppression. In this case, without CoT-based cold-start training, the model tends to converge to a risk-averse strategy: simply outputting answers without generating reasoning, so as to maintain stability. When AMTR is applied but the thinking mask is removed, additional constraints hinder the diversity of reasoning generation, thereby degrading model performance, particularly in reasoning-intensive scenarios (e.g., the RSVG dataset). In contrast, A2GRPO encourages constructive reasoning while removing unintended constraints on the thinking process, thereby expanding the reasoning exploration space. As a result, our method effectively activates reasoning while maintaining high answer accuracy, ultimately enhancing the model’s emergent reasoning capability.
@@
In summary, the results in Tables VIVIII demonstrate that, by leveraging hard samples and proper RL strategies (i.e., AMTR and the thinking mask), we successfully trigger the emergence of thinking in MLLMs on geospatial scenes, a capability that does not appear after SFT.
IV-D Visualization
@@
(a) Relationship between thinking and answering, where the statistics are computed from GeoZero’s prediction results on the DIOR-RSVG test set. (Best view in zoom.)
@@
Thinking v.s. Answering. To further investigate how reasoning relates to final predictions, we analyze the relationship between answer accuracy and both reasoning length and reasoning quality score $s_{t}$. Specifically, all samples are grouped based on their reasoning length and quality score, and the average accuracy within each group is computed. As shown in Figure 4(a)(a), accuracy increases with longer reasoning chains, indicating that extended reasoning generally benefits answer correctness. However, when the reasoning becomes excessively long, i.e., overthinking, the performance slightly drops. Similarly, Figure 4(a)(b) reveals a positive correlation between $s_{t}$ and answer accuracy, indicating that samples with higher reasoning quality scores tend to achieve higher accuracy. Overall, these results suggest a strong association between reasoning quality and task performance, supporting our design choice of encouraging explicit reasoning during RL. More fine-grained analyses are provided in Section-I of the supplementary material.
@@ Appendix B Can General MLLMs Perform Geospatial Reasoning?
To measure whether the model performs explicit reasoning, we adopt a simple yet permissive criterion: a sample is considered as “reasoned” if there exist any texts appearing before the “<answer>” tag in the model’s output. Based on this rule, we calculate the proportion of samples exhibiting reasoning behavior (referred to as the ”thinking activation rate”), and summarize the results in Table IX.
TABLE IX: Accuracy and thinking activation rate (%) of different models on the UCM and RSVG datasets. GeoZero uses the system prompt (SP) by default. w/SP: with the system prompt (Type1) shown in Figure 8. Acc.: accuracy. TAR: thinking activation rate.
@@ Appendix C Instruction Formatting of GeoZero-Raw
We apply instruction formatting to the GeoZero-Raw dataset. To distinguish different tasks, following the conventions of existing remote sensing MLLMs [28, 68], we prepend task descriptors to user instructions, where [cls], [grounding], [vqa], and [caption] represent scene classification (SC), visual grounding (VG), visual question answering (VQA), and image captioning (IC) tasks, respectively. For samples that do not belong to these four vision–language tasks, no task descriptors are added. Furthermore, to help general-purpose MLLMs better understand the expected output format for different remote sensing vision–language tasks, we introduce task-specific textual hints in addition to the original user instructions.. For example, for VG tasks, the instruction is reformulated as: “Please provide the location of the object in the image: {original grounding text}.” To prevent the model from over-relying on these textual hints, we randomly decide whether to include them with a probability of 50%. In addition, to further enhance model generalization, we design more than 20 variations of textual hints for each task, from which one is randomly selected for each sample. The textual hints for different tasks are shown in Figure 20-23, while the converted samples are presented in Figure 24.
@@ Appendix G Task-Specific Answer Reward $r_{a}$
TABLE X: Training hyperparameters for evaluating on different datasets used in scene classification, visual grounding, visual question answering, image captioning tasks and comprehensive benchmarks with GeoZero (w/ RFT).
@@ Appendix H Reinforcement Fine-tuning Settings
The hyperparameters used for reinforcement fine-tuning on different datasets are summarized in Table X. All other settings follow the reinforcement learning stage described in the main paper.
@@ Appendix I Further Investigation of Reasoning Behavior
I-A Overall Characteristics of Model Reasoning.
Figure 20(a)(a)–(f) illustrates the distribution of average sample accuracy across different reasoning-related metrics. It should be noted that in the main paper, we conducted a preliminary analysis using two variables: the word-level reasoning length and the thinking-quality score $s_{t}$ (see (a) and (f)). Here, we take a closer look by further decomposing $s_{t}$ into its internal factors. Specifically, we additionally analyze four variables: the normalized length score $l_{s}(T)$, the non-redundancy ratio 1d(T)1-d(T) (which serves as a redundancy penalty in $q_{t}$), the structural thinking score $q_{t}$, and the semantic diversity bonus $b_{d}$. Furthermore, we examine how the sample frequency is distributed across different value intervals of each metric (see (g)–(l)), which helps reveal the overall distribution of reasoning characteristics within the test data. Next, we provide a detailed analysis of each metric individually.
@@
    Reasoning Length. Figure 20(a)(b) shows a positive correlation between $l_{s}(T)$ and accuracy: samples with more appropriate reasoning lengths tend to yield more reliable predictions, consistent with the findings in the main paper. Encouragingly, we find that for the vast majority of samples, the predicted reasoning lengths fall within the expected range, as shown in Figure 20(a)(h), where more than 80% of the samples achieve very high $l_{s}(T)$ values (greater than 0.8). Figure 20(a)(g) further reveals that most reasoning chains have a length between 40 and 80 words, which precisely corresponds to the full-score interval defined in Equation (8).
@@
    Non-Redundancy. As shown in Figures 20(a)(c) and (i), almost no accuracy bars appear when $1-d(T)\leq 0.5$, while the vast majority of samples have values above 0.85. This suggests that the model’s reasoning chains exhibit strong non-redundancy across most samples. However, Figure 20(a)(c) also indicates that variations in reasoning redundancy have negligible influence on model performance.
    (a) The distribution of average accuracy and sample frequency on thinking-related metrics, where the statistics are computed from GeoZero’s predictions on the 7500 samples from the DIOR-RSVG [67] test set. The accuracy of samples are represented by IoU score.
    (b) The distribution of thinking-related variables with respect to accuracy, where the statistics are computed from GeoZero’s predictions on 7,500 samples from the DIOR-RSVG [67] test set. The sample accuracy is represented by the IoU score. The intervals are determined based on the quantiles of the IoU distribution, resulting in non-uniform spacing between the bars.
@@
    Structural Thinking. Figure 20(a)(d) clearly demonstrates that samples with low structural thinking scores $q_{t}$ tend to have low average accuracy, whereas those with higher $q_{t}$ consistently achieve better performance. This indicates that a well-structured reasoning process is a fundamental requirement for maintaining reliable predictions. However, the figure also shows that further increasing $q_{t}$ does not lead to substantial performance gains, implying that, at this stage, the semantic content of reasoning (such as semantic diversity; see (e)) may play a more dominant role. Moreover, Figure 20(a)(j) shows that GeoZero produces well-structured reasoning chains for the majority of samples, reflecting its coherent reasoning capability.
@@
    Semantic Diversity. Figure 20(a)(e) further examines the semantic diversity bonus $b_{d}$. We observe a non-monotonic relationship: moderate semantic diversity correlates with improved accuracy, as it reflects coherent multi-step reasoning. Once $b_{d}$ becomes excessively large, however, the reasoning often degenerates into fragmented or even contradictory statements, suggesting that the model is no longer performing genuine reasoning, which in turn hurts accuracy. Fortunately, Figure 20(a)(k) shows that for the vast majority of samples, the generated reasoning chains have $b_{d}$ values below 0.45, where semantic diversity plays a positive and constructive role.
@@
    Overall Reasoning Quality. Finally, Figures 20(a)(f) and (l) jointly indicate that GeoZero produces reasoning chains with high thinking-quality scores ($s_{t}$) for the majority of samples, which underpins the model’s overall performance across the test set. We also observe that some samples achieve good average accuracy despite relatively low $s_{t}$ values. This phenomenon corresponds to the distinct leftmost bar in Figure 20(a)(j), where the structural thinking scores ($q_{t}$) are also low. This may be attributed to easier cases in which the model can make correct predictions without maintaining a highly organized reasoning structure, e.g., by exhibiting mild redundancy in expression, as seen in the third-to-last bar of Figure 20(a)(i).
@@
I-B Reasoning Behavior across Samples.
To further understand the reasoning behavior of GeoZero, we visualize the average values of reasoning metrics as a function of the sample prediction accuracy [^1] Alternatively, one could employ frontier proprietary models to assess sample difficulty. However, due to the high API costs, we instead use GeoZero’s own prediction accuracy as a practical proxy. (measured by IoU for grounding tasks), as shown in Figure 20(b). The bars are more densely distributed in the high-IoU regions, indicating that the test data is dominated by relatively easy samples. Nevertheless, a few highly challenging cases also exist, as evidenced by the rightmost bars corresponding to IoU values close to zero.
@@
Interestingly, for the easiest samples (IoU $\in$ [0.95, 1]), the model tends to produce even longer reasoning chains. This leads to mild redundancy, which slightly affects the reasoning-chain quality score $s_{t}$ through the structural thinking score $q_{t}$, as reflected by the lower values of the leftmost bars in Figure 20(b)(c), (d), and (f). However, since these samples are inherently simple, such redundancy does not compromise the model’s prediction accuracy.